OpenStudy (anonymous):

What is x? 3x+4y=70 and 8x-3y=9

OpenStudy (abb0t):

the \(23^{rd}\) letter in the alphabet..

OpenStudy (anonymous):

You can either do elimination or substitution. Which one are you familiar with

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

and more elimenation... its easier

OpenStudy (anonymous):

you should use elimination

OpenStudy (anonymous):

First, decide which variable you are going to get rid of. Which one do you want to eliminate?

OpenStudy (anonymous):

how 'bout y?

OpenStudy (anonymous):

Ok, so multiply the first equation by 3 and the second equation by 4

OpenStudy (anonymous):

Multiply equations by 3 & by 4 respectively, you'll end up with 9x+12y=210, 32x-12y= 36, then you'd be able to do what @jprulz said

OpenStudy (anonymous):

9x+12y=210 and 32x - 12y = 36 ?

OpenStudy (anonymous):

Yes, then add the two together and y will be gone.

OpenStudy (anonymous):

ok. den it is 41x=246?

OpenStudy (anonymous):

Solve for x now.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Yes, now plug that into one of the equations to solve for y

OpenStudy (anonymous):

ok thanks!!!!

OpenStudy (anonymous):

24+4y=70 4y=70-24 i get 11.5

hero (hero):

3x+ 4y = 70 8x - 3y = 9 Add them together: 11x + y = 79 Isolate y: y = 79 - 11x Substitute y into the first equation: 3x + 4(79 - 11x) = 70 3x + 316 - 44x = 70 Isolate x: 316 - 70 = 44x - 3x 246 = 41x 246/41 = x 6 = x Substitute back in and solve for y: 3x + 4y = 70 3(6) + 4y = 70 18 + 4y = 70 4y = 70 - 18 4y = 52 y = 52/4 y = 13

hero (hero):

@britanny_

OpenStudy (anonymous):

yesh?

OpenStudy (anonymous):

oh ok that makes sense now !!!!!!!!!!!!!! Thanks!!! Thanks!!! Thanks!!! Thanks!!! Thanks!!!